# Survey in Characterization of Semantic Change

### v4

@@ 4.4.3 Orientation
Methods based on word embeddings have been widely used for many sentiment analysis tasks (Raffel2019ExploringTL) and to investigate sentiment at the lexicon level (Ito2020WordLevelCS). In the context of LSC characterization, Fonteyn2021AdjustingSA used the distance between the word vectors of ‘good’ and ‘bad’ to measure the polarity of the expression ‘to death,’ as illustrated in Figure 8.
Similarly, Hellrich2018ModelingWE Xie2019TextbasedIO employed diachronic static embeddings to analyze the sentiment of words over time. Hellrich et al. train word embeddings in COHA from 1810-1830 as $C_{1}$ and a VAD lexicon Warriner2013NormsOV as $\mathfrak{R}(C_{2})$, tracking for orientation of words the valence-arousal-dominance framework. They access the polarity of the word based on the k-nearest-neighbors from a VAD sentiment lexicon. Xie et al. use ‘moral’ seed words to compare, like ‘cheating’ and ’fairness,’ also for COHA as shown in Figure 9.
Figure 8: Adaptation from Fonteyn2021AdjustingSA. The line plot shows the evolution of polarity of the multi-word ‘to death.’ It went from a negative concept to a more positive one, ameliorating the dominant sense.
Figure 9: Illutration adapted from Xie2019TextbasedIO. The figure shows how moral sentiment toward slavery, democracy, and gay evolved over two centuries, mapped in a 2D embedding space.
@@ 4.5 Synthesis and Summary
Word frequency-based methods are simple to implement, offer an easily interpretable analysis, and require a relatively small corpus of data. However, these methods struggle with capturing word similarity (Mikolov2013EfficientEO) and compositionality (Mitchell2010CompositionID), limiting their semantic analysis capabilities. While N-grams slightly improve compositionality, they are constrained by Zipf’s law (Zipf1949HumanBA), requiring an exponential increase in data to increase the size of N. Additionally, frequency alone cannot distinguish between sense broadening and narrowing, requiring extra steps to handle polysemy. Co-occurrence tables also fail to correlate well with polysemy or semantic shifts. In parallel studies, these methods perform poorly in detecting word orientation (Socher2013RecursiveDM) and struggle with contexts involving sarcasm, irony, or negation due to their limited contextual awareness. Topic modeling is a powerful text analysis tool with many good libraries implementing state-of-the-art algorithms, and it doesn’t require supervision or data annotation. The downside is that these methods require manual interpretation, as statistically derived word clusters can be ambiguous. Automated topic selection using dictionary-based filtering aids in neologism detection but remains insufficient for change characterization. Metrics like coherence and diversity, along with visualization techniques, help reduce manual effort.
@@
Graph-based approaches rely heavily on annotated and structured data, which is often costly or, in some domains, unfeasible. However, Large Language Models enable automatic annotation and structuring of information. We have yet to fully leverage modern structures like knowledge graphs for LSC detection. Resources like Wikidata could help model semantic shifts (Giulianelli2023InterpretableWS; Tang2023CanWS), such as metonymization. Existing lexical databases (WordNet, BabelNet, ConceptNet) assist in change detection but lack temporal information, limiting their effectiveness.
@@
Word embeddings often exhibit instability, where minor corpus variations alter vector representations, especially for low-frequency or polysemous words (Hellrich2016Bad; Gladkova2016Intrinsic; Gnther2019Vector). Solutions such as larger embeddings, quantization, and alignment help mitigate this instability (Hamilton2016DiachronicWE); however, the meaning conflation deficiency (CamachoCollados2018FromWT) remains an issue, as a single vector—even when contextualized—does not provide enough information to differentiate fine-grained senses (Ballout2024FOOLMI; Periti2024AnalyzingSC). Subsequent work has explored various modifications to transformer architectures to enable better representation of meaning (Zhang2021DisentanglingRO; Tay2022UL2UL), but these approaches remain under-explored in the semantic change literature.
@@
Contextualized embeddings (e.g., from ELMo, BERT) enhance sense detection but can produce overly contextualized vectors, resulting in unreliable diachronic analysis and false positives (Kutuzov2022ContextualizedEF). Recent methods using contrastive learning and similarity-based training enhance isotropy and sense differentiation (Gao2021SimCSESC; Reimers2019SentenceBERTSE). Other approaches, like density-based embeddings (e.g., word2gauss), aim to model polysemy directly. The current state of the art in LSC identification fine-tunes cross-lingual models trained on Word Sense Disambiguation (WSD) tasks, benefiting from transfer learning across languages.
@@
While generative Artificial Intelligence (AI), particularly LLMs, is rapidly transforming NLP, its application to the systematic characterization of semantic change in diachronic data remains largely unexplored. However, some initial studies have successfully leveraged these models for related tasks, such as generating synthetic data for LSC detection (Cassotti2024Using) and automatically comparing word senses in context (Sa2024SemanticCC).
@@
biemann2006chinese X - - Graph
Sagi2009SemanticDA X - - Frequency
Lau2012WordSI X - - Topics
Cook2013ALA X - - Topics
Tang2013SemanticCC X X - Frequency
Lau2014LearningWS X - - Topics
Mitra2015AnAA X - - Graph
Emms2016DynamicGM X - - Frequency
Mathew2017Adapting X - - Graph and Topics
Tahmasebi2017FindingIW X - - Graph
Ryskina2020WhereNW X - - Embeddings
Giulianelli2020AnalysingLS X - - Embeddings
moss2020detecting X - - Embeddings
Ehmller2020SenseTD X - - Graph
Yksel2021SemanticCD X - - Embeddings
Inoue2022InfiniteSA X - - Topics
Bochkarev2022NeuralNA X - - Frequency
Lautenschlager2024DetectionON X - - Embeddings
Schlechtweg2017GermanIF - X - Frequency
Fonteyn2021AdjustingSA - X X Embeddings
Maudslay2022MetaphoricalPD - X - Embeddings
Cook2010AutomaticallyIC - - X Frequency
Jatowt2014AFF - - X Frequency
Buechel2016FeelingsFT - - X Frequency
Hamilton2016InducingDS - - X Graph
Hellrich2018ModelingWE - - X Embeddings
Xie2019TextbasedIO - - X Embeddings
@@
Dimension
(Broadening/
Narrowing)
Frequency
@@
Relation
(Metaphor/
Metonymy)
Frequency
@@
Orientation
(Amelioration/
Pejoration)
Frequency
@@ 5 Formalization
To the core of this formalization, a central challenge arises in defining what ”meaning” is. While linguistic literature often invokes terms like ‘concepts’ and ‘sentiments’ to describe meaning (koch2016meaning; traugott2017semantic; Blank2003Polysemy), these notions are rarely operationalized in a way suitable for formal modeling. This lack of clarity makes the computational implementation of semantic analysis particularly complex.
To address this, we adopt a minimal form of sense objectivism; the idea that word senses can be treated as identifiable units at specific points in time. Where the senses were conveyed by a community of speakers during the production of the corpus under study. While absolute objectivity is philosophically untenable (Geeraerts1997DiachronicPS; wittgenstein2009philosophical, pg. 182, pg. 32), this abstraction is both methodologically necessary and practically effective. It mirrors other linguistic constructs like phonemes: not directly observable or in practical terms precise, yet essential for systematic analysis. Indeed, the very concept of semantic change presupposes a stable referent; without identifiable senses, we cannot speak coherently of change; change compared to what?
@@
We build our formalization by posing the meaning of a word as a set of senses, i.e., glosses, inspired by HockJoseph2019Semantic, and define all constructions in set theory. First, we assume that for a set of text corpora $T$, there is a universe of senses $S_{T}$ that holds all senses for these corpora. Given an alphabet $\alpha$, we define $\mathcal{S}$ as a function that takes a string from the Kleene closure of $\alpha$, where $\alpha^{*}=V$ (the possible vocabulary), and returns a subset of $S_{T}$. For instance, given a word $w\in V$ and a corpus $t\in T$, the subset of senses $\mathcal{S}(w,t)$ for the universe of senses ($S_{t}$) of the word $w$ in this corpus is:
@@
in simple terms, there are objective senses of the word ‘a’ in corpus ‘x’ that do not appear in corpus ‘y’ for the word ‘b,’ or vice versa. This underlying assumption is made implicitly in most previous works, even though it is rarely stated explicitly, e.g., when Hamilton2016DiachronicWE and Inoue2022InfiniteSA state that the word ‘record’ has a different meaning because it gained the sense of ‘tape/music’ when comparing the word in corpora before and after 1920.
@@ 5.1 Semantic Change
koch2016meaning presents a definition of meaning change that combines semasiological and onomasiological perspectives. In Figure 10, ‘belly’ was initially associated with the source concept ‘bag’ (M1). Later, it acquired the meaning of the target concept ‘body part between the breast and the thighs’ (M2). In computational linguistics, we primarily focus on the semasiological view: our interest lies in the word itself and the evolution of its meanings, rather than the origins of the concepts it represents.
Figure 10: Definition of semantic change adapted from koch2016meaning.
@@ 5.2 Change in Dimension
This definition is compatible with the concept of innovative/reductive meaning change from koch2016meaning, in which words acquire new senses (sememes) by being used with different meanings. It differs, however, from the more general notion of broadening, where words may also lose intension (the precision of their meaning), making them vaguer.
@@
In this survey, we define broadening as the degree of polysemy in a word. This is aligned with the NLP literature, given its application in data analysis and its measurable nature, compared to the notion of intension*. We differ from the definition in Tang2013SemanticCC, where broadening is a superclass of metaphorization and metonymization, as the literature reviewed here does not perform this aggregation.
@@ 5.3 Change in Relation
Metaphorization or metonymization occurs when a word takes on a new meaning that inherits qualities from its original meaning through a figurative relationship the speaker aims to convey (lakoff2008metaphors, p. 3-6). The original meaning is invoked by establishing a material or abstract relation with the initial sense, where the listener is able to resolve the non-literal target concept Fass1988MetonymyAM. Material relations are built on physical-world contiguity (e.g., replacing an institution with its location, as in “Washington decided…”). In contrast, abstract relations are formed through conceptual similarity (e.g., applying human body parts to objects, as in “the arm of the chair”).
The theory defining metaphor and metonymy is often informal, relying on subjective perception or overlapping definitions. For example, Choi2021Melbert defines metaphor as a case where the “literal meaning of a word is different from its contextual meaning,” which also applies to metonymy. As discussed by koch2016meaning, there is no precise rule to define a metonymy. Some computational work, such as Schlechtweg2017GermanIF, measures a word’s entropy to capture metaphorization, but it is not clear that this measure can differentiate a metaphor from a metonymy or a completely new, unrelated sense (e.g., ‘Apple’ in a technology context).
@@
$\mathcal{R}(w,t)=\left\{\begin{aligned} \sum^{N-1}_{i=1}\sum^{N}_{j=i+1}{\mathit{l}(s_{i},s_{j})}\hskip 30.00005pts_{i,j}\in\mathcal{S}(w,t)\\ 0\hskip 60.00009ptN\leq 1,\end{aligned}\right$. (6)
@@ 5.4 Change in Orientation
A word’s change in orientation (positive or negative connotation) reflects amelioration or pejoration in its usage from one corpus to another. This is measured by analyzing how the collective perception of a word’s meaning evolves (campbell2013historical, p. 227-229). A classic example is ‘terrific,’ which originally meant ‘terrible’ but shifted to mean ‘formidable’ or ‘impressive.’
While some frameworks like Buechel2016FeelingsFT use a multi-dimensional Valence-Arousal-Dominance setting, most sentiment analysis research focuses on a simpler polarity setting (Cui2023SurveyOS). Therefore, we restrict our formalism to this more general positive/negative view to align with current research and applications (Susanto2020TheHM).
@@ 5.5 Semantic Change Identification
Core to computational approaches, words in a corpus are not compared directly to one another; instead, they are first represented using one of the many methods reviewed in this survey. A common step taken by all authors is to represent the word ‘a‘ in corpus ‘x‘, denoted as $\mathfrak{R}(a,x)$, and then compare this representation to $\mathfrak{R}(b,y)$. For example, moss2020detecting represents the word ‘a‘ as a Gaussian embedding obtained after training a language model on corpus ‘x‘. Similarly, Ehmller2020SenseTD represents ‘a‘ using a co-occurrence graph derived from corpus ‘x‘. Meanwhile, Inoue2022InfiniteSA partitions the original corpus into 20-year intervals and, for each resulting corpus ‘x‘, obtains a distribution of topics for the word ‘a‘.
@@
Schlechtweg2020SimulatingLS proposed two notions of semantic change: binary classification, where $y\in\{0,1\}$, and graded (continuous) semantic change, where $y\in\mathbf{R}$ is a distance between word representations [^3] While Schlechtweg2020SimulatingLS considered graded change for a particular purpose (i.e., Jensen-Shannon distance), we generalize $y$ to any applicable distance measure..
The binary classification task aligns with classic LSC literature (blank2012prinzipien, p. 113) and human perception, which is often influenced by sensory thresholds (smith2008biology, p. 34). This view is interested in the gain or loss of senses. There is no general threshold criterion for the binary setting; it’s not enough for representations to be different, the magnitude of this difference should align with human judgment.

### v5

@@ 4.4.3 Orientation
Methods based on word embeddings have been widely used for many sentiment analysis tasks (Raffel et al., 2019) and to investigate sentiment at the lexicon level (Ito et al., 2020). In the context of LSC characterization, Fonteyn and Manjavacas (2021) used the distance between the word vectors of ‘good’ and ‘bad’ to measure the polarity of the expression ‘to death,’ as illustrated in Figure 8.
Similarly, Hellrich et al. (2018) Xie et al. (2019) employed diachronic static embeddings to analyze the sentiment of words over time. Hellrich et al. train word embeddings in COHA from 1810-1830 as $C_{1}$ and a VAD lexicon Warriner et al. (2013) as $\mathfrak{R}(C_{2})$, tracking for orientation of words the valence-arousal-dominance framework. They access the polarity of the word based on the k-nearest-neighbors from a VAD sentiment lexicon. Xie et al. use ‘moral’ seed words to compare, like ‘cheating’ and ’fairness,’ also for COHA as shown in Figure 9.
Figure 8: Adaptation from Fonteyn and Manjavacas (2021). The line plot shows the evolution of polarity of the multi-word ‘to death.’ It went from a negative concept to a more positive one, ameliorating the dominant sense.
Figure 9: Illutration adapted from Xie et al. (2019). The figure shows how moral sentiment toward slavery, democracy, and gay evolved over two centuries, mapped in a 2D embedding space.
@@ 4.5 Synthesis and Summary
Word frequency-based methods are simple to implement, offer an easily interpretable analysis, and require a relatively small corpus of data. However, these methods struggle with capturing word similarity (Mikolov et al., 2013) and compositionality (Mitchell and Lapata, 2010), limiting their semantic analysis capabilities. While N-grams slightly improve compositionality, they are constrained by Zipf’s law (Zipf, 1949), requiring an exponential increase in data to increase the size of N. Additionally, frequency alone cannot distinguish between sense broadening and narrowing, requiring extra steps to handle polysemy. Co-occurrence tables also fail to correlate well with polysemy or semantic shifts. In parallel studies, these methods perform poorly in detecting word orientation (Socher et al., 2013) and struggle with contexts involving sarcasm, irony, or negation due to their limited contextual awareness. Topic modeling is a powerful text analysis tool with many good libraries implementing state-of-the-art algorithms, and it doesn’t require supervision or data annotation. The downside is that these methods require manual interpretation, as statistically derived word clusters can be ambiguous. Automated topic selection using dictionary-based filtering aids in neologism detection but remains insufficient for change characterization. Metrics like coherence and diversity, along with visualization techniques, help reduce manual effort.
@@
Graph-based approaches rely heavily on annotated and structured data, which is often costly or, in some domains, unfeasible. However, Large Language Models enable automatic annotation and structuring of information. We have yet to fully leverage modern structures like knowledge graphs for LSC detection. Resources like Wikidata could help model semantic shifts (Giulianelli et al., 2023; Tang et al., 2023), such as metonymization. Existing lexical databases (WordNet, BabelNet, ConceptNet) assist in change detection but lack temporal information, limiting their effectiveness.
@@
Word embeddings often exhibit instability, where minor corpus variations alter vector representations, especially for low-frequency or polysemous words (Hellrich and Hahn, 2016; Gladkova and Drozd, 2016; Günther et al., 2019). Solutions such as larger embeddings, quantization, and alignment help mitigate this instability (Hamilton et al., 2016b); however, the meaning conflation deficiency (Camacho-Collados and Pilehvar, 2018) remains an issue, as a single vector—even when contextualized—does not provide enough information to differentiate fine-grained senses (Ballout et al., 2024; Periti et al., 2024). Subsequent work has explored various modifications to transformer architectures to enable better representation of meaning (Zhang et al., 2021; Tay et al., 2022), but these approaches remain under-explored in the semantic change literature.
@@
Contextualized embeddings (e.g., from ELMo, BERT) enhance sense detection but can produce overly contextualized vectors, resulting in unreliable diachronic analysis and false positives (Kutuzov et al., 2022). Recent methods using contrastive learning and similarity-based training enhance isotropy and sense differentiation (Gao et al., 2021; Reimers and Gurevych, 2019). Other approaches, like density-based embeddings (e.g., word2gauss), aim to model polysemy directly. The current state of the art in LSC identification fine-tunes cross-lingual models trained on Word Sense Disambiguation (WSD) tasks, benefiting from transfer learning across languages.
@@
While generative Artificial Intelligence (AI), particularly LLMs, is rapidly transforming NLP, its application to the systematic characterization of semantic change in diachronic data remains largely unexplored. However, some initial studies have successfully leveraged these models for related tasks, such as generating synthetic data for LSC detection (Cassotti et al., 2024) and automatically comparing word senses in context (de Sá et al., 2024).
@@
Biemann (2006) X - - Graph
Sagi et al. (2009) X - - Frequency
Lau et al. (2012) X - - Topics
Cook et al. (2013) X - - Topics
Tang et al. (2013) X X - Frequency
Lau et al. (2014) X - - Topics
Mitra et al. (2015) X - - Graph
Emms and Jayapal (2016) X - - Frequency
Mathew et al. (2017) X - - Graph and Topics
Tahmasebi and Risse (2017) X - - Graph
Ryskina et al. (2020) X - - Embeddings
Giulianelli et al. (2020) X - - Embeddings
Moss (2020) X - - Embeddings
Ehmüller et al. (2020) X - - Graph
Yüksel et al. (2021) X - - Embeddings
Inoue et al. (2022) X - - Topics
Bochkarev et al. (2022) X - - Frequency
Lautenschlager et al. (2024) X - - Embeddings
Schlechtweg et al. (2017) - X - Frequency
Fonteyn and Manjavacas (2021) - X X Embeddings
Maudslay and Teufel (2022) - X - Embeddings
Cook and Stevenson (2010) - - X Frequency
Jatowt and Duh (2014) - - X Frequency
Buechel et al. (2016) - - X Frequency
Hamilton et al. (2016a) - - X Graph
Hellrich et al. (2018) - - X Embeddings
Xie et al. (2019) - - X Embeddings
@@
Dimension (Broadening/ Narrowing) Frequency
@@
Relation (Metaphor/ Metonymy) Frequency
@@
Orientation (Amelioration/ Pejoration) Frequency
@@ 5 Formalization
To the core of this formalization, a central challenge arises in defining what ”meaning” is. While linguistic literature often invokes terms like ‘concepts’ and ‘sentiments’ to describe meaning (Koch, 2016; Traugott, 2017; Blank and others, 2003), these notions are rarely operationalized in a way suitable for formal modeling. This lack of clarity makes the computational implementation of semantic analysis particularly complex.
To address this, we adopt a minimal form of sense objectivism; the idea that word senses can be treated as identifiable units at specific points in time. Where the senses were conveyed by a community of speakers during the production of the corpus under study. While absolute objectivity is philosophically untenable (Geeraerts, 1997; Wittgenstein, 2009, pg. 182, pg. 32), this abstraction is both methodologically necessary and practically effective (as we target static corpora). It mirrors other linguistic constructs like phonemes: not directly observable or in practical terms precise, yet essential for systematic analysis. Indeed, the very concept of semantic change presupposes a stable referent; without identifiable senses, we cannot speak coherently of change; change compared to what?
@@
We build our formalization by posing the meaning of a word as a set of senses, i.e., glosses, inspired by Hock and Joseph (2019, p. 194), and define all constructions in set theory. First, we assume that for a set of text corpora $T$, there is a universe of senses $S_{T}$ that holds all senses for these corpora. Given an alphabet $\alpha$, we define $\mathcal{S}$ as a function that takes a string from the Kleene closure of $\alpha$, where $\alpha^{*}=V$ (the possible vocabulary), and returns a subset of $S_{T}$. For instance, given a word $w\in V$ and a corpus $t\in T$, the subset of senses $\mathcal{S}(w,t)$ for the universe of senses ($S_{t}$) of the word $w$ in this corpus is:
@@
in simple terms, there are objective senses of the word ‘a’ in corpus ‘x’ that do not appear in corpus ‘y’ for the word ‘b,’ or vice versa. This underlying assumption is made implicitly in most previous works, even though it is rarely stated explicitly, e.g., when Hamilton et al. (2016b) and Inoue et al. (2022) state that the word ‘record’ has a different meaning because it gained the sense of ‘tape/music’ when comparing the word in corpora before and after 1920.
@@ 5.1 Semantic Change
Koch (2016, p. 23) presents a definition of meaning change that combines semasiological and onomasiological perspectives. In Figure 10, ‘belly’ was initially associated with the source concept ‘bag’ (M1). Later, it acquired the meaning of the target concept ‘body part between the breast and the thighs’ (M2). In computational linguistics, we primarily focus on the semasiological view: our interest lies in the word itself and the evolution of its meanings, rather than the origins of the concepts it represents.
Figure 10: Definition of semantic change adapted from Koch (2016, p. 23,25).
@@ 5.2 Change in Dimension
This definition is compatible with the concept of innovative/reductive meaning change from Koch (2016, p. 24-26), in which words acquire new senses (sememes) by being used with different meanings. It differs, however, from the more general notion of broadening, where words may also lose intension (the precision of their meaning), making them vaguer.
@@
In this survey, we define broadening as the degree of polysemy in a word. This is aligned with the NLP literature, given its application in data analysis and its measurable nature, compared to the notion of intension. We differ from the definition in Tang et al. (2013), where broadening is a superclass of metaphorization and metonymization, as the literature reviewed here does not perform this aggregation.
@@ 5.3 Change in Relation
Metaphorization or metonymization occurs when a word takes on a new meaning that inherits qualities from its original meaning through a figurative relationship the speaker aims to convey (Lakoff and Johnson, 2008, p. 3-6). The original meaning is invoked by establishing a material or abstract relation with the initial sense, where the listener is able to resolve the non-literal target concept Fass (1988). Material relations are built on physical-world contiguity (e.g., replacing an institution with its location, as in “Washington decided…”). In contrast, abstract relations are formed through conceptual similarity (e.g., applying human body parts to objects, as in “the arm of the chair”).
The theory defining metaphor and metonymy is often informal, relying on subjective perception or overlapping definitions. For example, Choi et al. (2021) defines metaphor as a case where the “literal meaning of a word is different from its contextual meaning,” which also applies to metonymy. As discussed by Koch (2016, p. 42-51), there is no precise rule to define a metonymy. Some computational work, such as Schlechtweg et al. (2017), measures a word’s entropy to capture metaphorization, but it is not clear that this measure can differentiate a metaphor from a metonymy or a completely new, unrelated sense (e.g., ‘Apple’ in a technology context).
@@
$\mathcal{R}(w,t)=\left\{\begin{aligned} \sum^{N-1}_{i=1}\sum^{N}_{j=i+1}{\mathit{l}(s_{i},s_{j})}\hskip 30.00005pts_{i},s_{j}\in\mathcal{S}(w,t)\\ 0\hskip 60.00009ptN\leq 1,\end{aligned}\right$. (6)
@@ 5.4 Change in Orientation
A word’s change in orientation (positive or negative connotation) reflects amelioration or pejoration in its usage from one corpus to another. This is measured by analyzing how the collective perception of a word’s meaning evolves (Campbell, 2013, p. 227-229). A classic example is ‘terrific,’ which originally meant ‘terrible’ but shifted to mean ‘formidable’ or ‘impressive.’
While some frameworks like Buechel et al. (2016) use a multi-dimensional Valence-Arousal-Dominance setting, most sentiment analysis research focuses on a simpler polarity setting (Cui et al., 2023). Therefore, we restrict our formalism to this more general positive/negative view to align with current research and applications (Susanto et al., 2020).
@@ 5.5 Semantic Change Identification
Core to computational approaches, words in a corpus are not compared directly to one another; instead, they are first represented using one of the many methods reviewed in this survey. A common step taken by all authors is to represent the word ‘a‘ in corpus ‘x‘, denoted as $\mathfrak{R}(a,x)$, and then compare this representation to $\mathfrak{R}(b,y)$. For example, Moss (2020) represents the word ‘a‘ as a Gaussian embedding obtained after training a language model on corpus ‘x‘. Similarly, Ehmüller et al. (2020) represents ‘a‘ using a co-occurrence graph derived from corpus ‘x‘. Meanwhile, Inoue et al. (2022) partitions the original corpus into 20-year intervals and, for each resulting corpus ‘x‘, obtains a distribution of topics for the word ‘a‘.
@@
Schlechtweg and im Walde (2020) proposed two notions of semantic change: binary classification, where $y\in\{0,1\}$, and graded (continuous) semantic change, where $y\in\mathbf{R}$ is a distance between word representations [^3] While Schlechtweg and im Walde (2020) considered graded change for a particular purpose (i.e., Jensen-Shannon distance), we generalize $y$ to any applicable distance measure..
The binary classification task aligns with classic LSC literature (Blank, 2012, p. 113) and human perception, which is often influenced by sensory thresholds (Smith, 2008, p. 34). This view is interested in the gain or loss of senses. There is no general threshold criterion for the binary setting; it’s not enough for representations to be different, the magnitude of this difference should align with human judgment.